# Peer-Aware Cost Estimation in Nonlinear General-Sum Dynamic Games for Mutual Learning and Intent Inference

### v1

@@ IV-A Case Study 1: Assistive Lunar Lander
Figure 2: Results of the assistive lunar lander example. (a) Trajectory comparison between the complete information game, N-PACE, and the expert peer assumption. (b) and (c) show each agent’s learning performance for the final desired position under different algorithms.
IV-B Case Study 2: Autonomous Vehicles in Lane Merging
Figure 3: Trajectories of the lane-merging case study. (a) Under complete information, safe lane merging is ensured as the red car yields to the more aggressive blue car. (b) Under incomplete information, where each agent assumes the other is an expert during intent inference, a collision occurs. (c) N-PACE results show that when agents infer each other’s aggressiveness while also accounting for their peer’s learning, safe lane merging is achieved, with the red car yielding to the blue car.
[…]
$\dot{x}_{1}$ $=v_{1},\quad\dot{v}_{1}=a_{1},\quad\dot{x}_{2}=v_{2}\cos{\theta_{2}},\quad\dot{y}_{2}=v_{2}\sin{\theta_{2}},$
$\dot{\theta}_{2}$ $=\delta_{2},\quad\dot{v}_{2}=a_{2},\quad g_{blue}=0.1(x_{1}-25)^{2}+0.1v_{1}^{2}$
$+$ $\theta_{1}e^{-(d(t)-d_{\text{safe}})^{2}}+a_{1}^{2},\ g_{red}=1y_{2}^{2}+10\theta_{2}^{2}+0.1v_{2}^{2}$
$+$ $0.1(x_{2}-25)^{2}+\theta_{2}e^{-(d(t)-d_{\text{safe}})^{2}}+a_{2}^{2}+\delta_{2}^{2}.$
where $d(t)=(x_{1}-x_{2})^{2}+y_{2}^{2}$ is the distance between the two cars, $d_{\text{safe}}$ is the safe distance they aim to maintain, and $\theta_{k}$ is a parameter in the range $[0,100]$ that represents the aggressiveness of each agent (lower $\theta_{k}$ corresponds to more aggressive behavior).
Figure 4: Comparing parameter learning of each agent in case study 2 under N-PACE and ”treating others as Expert” with complete information, highlighting the superior performance of N-PACE and its fast convergence.
Although treating a learning agent as an expert and attempting to estimate their aggressiveness may not converge to the true parameter value, this approach does not always necessarily lead to a collision. In this example, we consider a specific scenario and initial conditions where the complete information game ensures safe lane merging, and our N-PACE algorithm also enables fast and accurate parameter learning; however, under the same initial conditions and learning dynamics, treating the peer agent as an expert leads to not only incorrect parameter learning but also collisions and safety issues. Each agent initially assumes that the other agent is a non-aggressive, safety-conscious driver with $\hat{\theta}^{k}_{0}=100$. However, both agents exhibit some level of aggressiveness: the first agent (the blue car) is highly aggressive with $\theta_{1}=5$, while the second agent (the red car) has a moderate aggressiveness level of $\theta_{2}=80$. As shown in Fig. 3, in the complete information game scenario, agent 2 (the red car) yields to agent 1 (the blue car), successfully avoiding a collision. However, in the incomplete information game, where agents treat their peers as experts, the biased learning process leads not only to incorrect parameter estimations but also to overlapping safety regions in their trajectories, resulting in safety risks. In contrast, N-PACE, due to its unbiased learning process, converges quickly to the true parameter values and avoids collisions. As a result, its behavior closely resembles that of the complete information game, with the red car yielding to the blue car, albeit with some latency.
IV-C Case Study 3: Autonomous Vehicles in an Interaction
Figure 5: Trajectory $(x,y)$ results of 300 simulations of the two-player incomplete information intersection game with random values of aggressiveness. Gray boxes indicate the collision area. (a) The complete information game results in no collisions. (b) The incomplete information game, where agents assume their peer is a complete information agent, results in 12 collision cases out of 300. (c) N-PACE results in only 4 collisions. (d) N-PACE combined with intent communication achieves 0 collisions and closest trajectories to those in the complete information game.
In our third case study, to extensively study N-PACE and the intent communication approach proposed in (7), we conduct Monte Carlo simulations of an intersection driving game between two agents under incomplete information, similar to the scenario shown in Fig. 1. In this case study, two autonomous vehicles attempt to navigate through an intersection while being unaware of each other’s aggressiveness parameter, $\theta_{k}$. The agents are only able to control their acceleration along with the direction of their movement $x,y$. The dynamic of the game and each agent cost function follow as:
$\dot{x}_{1}$ $=v_{1},\quad\dot{v}_{1}=a_{1},\quad\dot{y}_{2}=v_{2},\quad\dot{v}_{2}=a_{2},$
[…]
where $d(t)=x_{1}^{2}+y_{2}^{2}$ is the distance between two vehicles and $d_{safe}$ is the safe distance they try to maintain. The aggressiveness of each agent is also defined same as Case Study 2. The game continues for $5$ sec with a sampling time of $\Delta{T}=0.1$ sec, resulting in 50 samples.
In this example, each agent’s aggressiveness parameter, $\theta_{k}$, takes values in the range $[20,50]$, where a lower $\theta_{k}$ indicates a more aggressive agent. In our Monte Carlo study, we sampled 300 aggressiveness values uniformly from $[20,50]$ for each agent. Using these 300 sampled values, we ran the game 300 times under four different scenarios: 1) Complete information, 2) The expert peer agent assumption, 3) N-PACE, 4) N-PACE with intent communication, as proposed in Sec. III-E, for different intent communication weighting values $\eta=0.1,1,10$.
The trajectories from these 300 simulations for different scenarios are shown in Fig. 5, where we plot the $x$-position of the first agent against the $y$-position of the second agent. The gray box in the middle represents the collision area in the center of the intersection. As expected, under the complete information game scenario, no collisions were observed. However, in the scenario where each agent assumes its peer is a complete information/expert agent and attempts to infer their objective, we observed 12 collisions out of 300 simulations. In contrast, when agents used N-PACE to infer each other’s objectives, only 4 out of 300 trajectories resulted in collisions—a $66\%$ reduction compared to the baseline approach. It is important to note that having 4 collisions under N-PACE is not unexpected, as N-PACE also has a limited convergence speed during the interaction. This convergence speed depends on factors such as the chosen learning dynamics and the persistence of excitation in control signals. Interesting results were observed when we applied the proposed intent communication framework from Sec. III-E to this problem. In Fig. 5 (d), we present the trajectory results for $\eta=1$, where agents communicate their intent. Under this scenario, no collisions were observed in any of the 300 simulations. This was also the case for $\eta=10$, while only a single collision occurred for $\eta=0.1$. These results highlight the improvements achieved in both inference accuracy and safety assurance using the proposed intent communication framework built on top of N-PACE. Finally, Fig. 6 presents the log of the Mean Square learning Error (MSE) of both agents across all 300 Monte Carlo simulations. We use $\log(1+\text{MSE})$ for better visualization and comparison. The results clearly demonstrate the superiority of N-PACE over the approach where each agent treats the other as an expert, in terms of learning error. Additionally, they highlight how intent communication improves learning performance, particularly when the signaling is intensified using higher $\eta$.

### v2

@@ 3.4. N-PACE for Intent Communication
Figure 2. Results of Case Study 1, the assistive lunar lander. (a) Trajectory comparison between baselines and N-PACE. (b) and (c) show each agent’s learning performance under different algorithms (d) comparison of final landing position error. (e) Final sum of intent learning error. (f) and (g) cumulative cost comparisons under different approaches for each agent.
Lemma 0 (Monotone reduction of intent estimation error under intent communication).
Assuming a fixed $t$, ego agent $k$, state $s_{t}$, peer action $a_{t}^{-k}$, current beliefs $(\hat{\theta}^{k}_{t},\hat{\theta}^{-k}_{t})$, an admissible action domain $\mathcal{A}_{t}^{k}\!\subseteq\!\mathbb{R}^{m_{k}}$, a base objective $C_{t}^{k}(a^{k})\!=\!J^{k}(s_{t},a^{k},a_{t}^{-k})$, and any learning dynamics $h^{-k}_{\ell}$. For any $0\leq\eta_{1}<\eta_{2}$, let action $a_{t}^{k,\eta}$ be any solution of the optimization problem (7) at each step $t$,the parameter learning error $E_{t}^{k}\!\big(a_{t}^{k,\eta}\big)=\big\|\hat{\theta}^{k}_{t+1}(a_{t}^{k,\eta})-\theta^{k}\big\|^{2}$ satisfies $E_{t}^{k}\!\big(a_{t}^{k,\eta_{2}}\big)\ \leq\ E_{t}^{k}\!\big(a_{t}^{k,\eta_{1}}\big).$ In particular, $\eta\mapsto E_{t}^{k}(a_{t}^{k,\eta})$ is non-increasing.
Proof.
See Appendix A. ∎
Corollary 0 (Intent communication never hurts).
Let $a_{t}^{k,0}\in\arg\min_{\mathcal{A}_{t}^{k}}C_{t}^{k}$ be any no-teaching optimizer and, for $\eta\geq 0$, let $a_{t}^{k,\eta}\in\arg\min_{\mathcal{A}_{t}^{k}}\Phi_{t}^{k,\eta}$ with $\Phi_{t}^{k,\eta}:=C_{t}^{k}+\eta E_{t}^{k}$. Then for every $\eta>0$, $E_{t}^{k}\!\big(a_{t}^{k,\eta}\big)\ \leq\ E_{t}^{k}\!\big(a_{t}^{k,0}\big).$ Moreover, if $\arg\min_{\mathcal{A}_{t}^{k}}C_{t}^{k}\,\cap\,\arg\min_{\mathcal{A}_{t}^{k}}E_{t}^{k}=\varnothing$ (i,e. disjoint argmin sets), then there exists $\eta^{\star}>0$ such that for all $\eta>\eta^{\star}$, $E_{t}^{k}\!\big(a_{t}^{k,\eta}\big)\ <\ E_{t}^{k}\!\big(a_{t}^{k,0}\big).$
Proof.
The non-strict inequality is Lemma 6 with $\eta_{1}=0$ and $\eta_{2}=\eta$. For strictness, Define
$C_{\star}:=\min_{\mathcal{A}_{t}^{k}}C_{t}^{k},\quad E_{0}:=\min_{a\in\arg\min C_{t}^{k}}E_{t}^{k}(a),\quad E_{\star}:=\min_{\mathcal{A}_{t}^{k}}E_{t}^{k}(a).$
Disjointness implies $E_{\star}<E_{0}$. Pick $a_{E}\in\arg\min E_{t}^{k}$ and set $\Delta C:=C_{t}^{k}(a_{E})-C_{\star}\geq 0$, $\Delta E:=E_{0}-E_{\star}>0$. Then
$\Phi_{t}^{k,\eta}(a_{E})-\Phi_{t}^{k,\eta}(a_{t}^{k,0})=\Delta C-\eta\,\Delta E.$
Choosing $\eta^{\star}:=\Delta C/\Delta E$ makes this negative for all $\eta>\eta^{\star}$, hence any minimizer $a_{t}^{k,\eta}$ satisfies $E_{t}^{k}(a_{t}^{k,\eta})\leq E_{t}^{k}(a_{E})=E_{\star}<E_{0}=E_{t}^{k}(a_{t}^{k,0})$. ∎
We will show the effectiveness of this approximation for intent communication and empirical evidence validating Lemma 6 and Corollary 7 in Case Study 3 in Sec. 4.
4. Case Studies
In this section, we show that explicitly modeling a peer’s learning dynamics with N-PACE improves task completion (Case Study 1) and safety (Case Studies 2–3) compare to baseline methods. We also evaluate the robustness of N-PACE to misspecified initial beliefs about the peer’s parameter learning (Case Study 2), and present both an ablation on initialization error and the effect of intent communication (Case Study 3). Across all case studies, the learning updates use gradient-descent-based–based Bayesian inference with a Gaussian belief. (Simulation Codes are available Github. mathematical details of baselines are available in the Appendix)
4.1. Case Study 1: Assistive Lunar Lander
Our first example focuses on the importance of N-PACE in task completion. In this example, we study a shared control problem, an assistive lunar lander scenario in which a human agent controls only the angle of the lunar lander $\phi$ by applying the torque $T$, while an autonomous agent controls the thrust force $F$ along the direction of the lunar lander’s angle (assuming no gravity). Both agents cooperate to land the lunar lander at the final position $(x_{f},y_{f})$, but with asymmetric information: the human knows only the final desired $x$-position $x_{f}$, while the autonomous agent knows only the final desired $y$-position $y_{f}$. As a result, the human must learn the parameter $\hat{y}_{f}$, and the autonomous agent must learn $\hat{x}_{f}$ based on observed interactions. The underlying nonlinear dynamics follow $\dot{x}=v_{x}$, $\dot{y}=v_{y}$, $\dot{\phi}=\omega$, $\dot{v}_{x}=F\sin{\phi}$, $\dot{\omega}=T$ the running costs for the human $g_{h}$ and the agent $g_{a}$ are defined as $g_{a}=10(x-\hat{x}_{f})^{2}+10(y-y_{f})^{2}+10(\phi-\frac{\pi}{2})^{2}+10(v_{x}^{2}+v_{y}^{2}+\omega^{2})+F^{2}$ and $g_{h}=10(x-x_{f})^{2}+10(y-\hat{y}_{f})^{2}+10(\phi-\frac{\pi}{2})^{2}+10(v_{x}^{2}+v_{y}^{2}+\omega^{2})+T^{2}.$ The game has a time horizon of $T=5s$ with a sampling time of $\Delta T=0.1s$ that is used to discretize the above dynamics. In this example, we set $x_{f}=5$ and $y_{f}=-5$, while the lunar lander starts at the initial position (5,8)(-5,8) with a zero initial angle. We conducted three simulations for this case study: 1) a complete information game where both agents have full knowledge of $x_{f}$ and $y_{f}$ as a ground truth baseline, 2) a game theoretic scenario where each agent treats its peer as an expert, assuming the other agent has complete information, 3)a non-game-theoretic model predictive control (MPC) inspired by the baselines in peters2024contingency where each agent uses a history of observed data and fits a linear policy to its partner policy parameters (including the unknown goal position) and using this policy perform a receding horizon MPC, and finally 4) a scenario in which agents operate according to N-PACE.
Figure 3. Trajectories of the lane-merging case study, the trajectories are plotted relative to the blue car coordinate (a) Safe lane merging under complete information (b) baseline methods, a risky behavior in merging from the aggressive driver (red). (c)Non-risky lane merging in N-PACE similar to the complete game, despite a 30% error in modeling the peer’s initial estimate.
Figure 4. a) & b) Robustness of N-PACE in intent inference under mismatched assumptions about the peer’s initial belief. c) & d) robustness to wrong choice of initial variance in peer’s GD-based Bayesian inference. Envelopes representing 50 mismatched runs.
Figure 2 shows a summary of the results, where Fig. 2 (a) presents the trajectory of the lunar lander under different algorithms, explicitly demonstrating the superiority of N-PACE compared to other approaches. It can be seen that the final position of the lunar lander in N-PACE is closer to the desired final position, and the entire trajectory aligns more closely with the complete information game trajectory (the local Nash equilibrium). In contrast, treating the other agent as an expert or the MPC-based baseline results in the lunar lander landing at a different position, emphasizing the importance of modeling the learning dynamics of the other agent for successful task completion. Figs. 2 (b) and (c) present the results of parameter learning performance for each agent under different algorithms, illustrating how N-PACE fast and successfully converges to the true final position goals in comparison to other baselines. Notably, treating the other agent as an expert can lead to a significant overshoot in the learning process. Intuitively, this may be due to a high initial bias in assuming that the peer agent has complete information. 2 (f) and (g) interestingly show how N-PACE achieves a lower cumulative cost for each agent compared to other methods in this case study.
4.2. Case Study 2: Lane Merging
This example aims to study the N-PACE robustness to imperfect modeling of the peers’ learning dynamics. In our second case study, we consider a lane-merging task between two autonomous vehicles. The first car (blue) moves in a straight line, while the second car (red) attempts to merge into its lane. The blue car can only accelerate or decelerate $a_{1}$, whereas the red car can control both its acceleration along its direction of motion $a_{2}$ and its steering angle velocity $\delta$. The game has a 5s horizon with a sampling time of $\Delta{T}=0.1s$ for dynamic discretization. During the game, Each agent is unaware of the aggressiveness level of others in their cost function and must learn it in real-time. The underlying dynamic is defined as $\dot{x}_{1}=v_{1}$, $\dot{v}_{1}=a_{1}$, $\dot{x}_{2}=v_{2}\cos{\phi_{2}}$, $\dot{y}_{2}=v_{2}\sin{\phi_{2}}$, $\dot{\phi}_{2}=\delta_{2},$ $\dot{v}_{2}=a_{2}$, and the running cost $g$ for each car is $g_{blue}=0.1(x_{1}-25)^{2}+0.1v_{1}^{2}+\theta_{1}e^{-(d(t)-d_{\text{safe}})^{2}}+a_{1}^{2},$ and $g_{red}=1y_{2}^{2}+10\phi_{2}^{2}+0.1v_{2}^{2}+0.1(x_{2}-25)^{2}+\theta_{2}e^{-(d(t)-d_{\text{safe}})^{2}}+a_{2}^{2}+\delta_{2}^{2}$. where $d(t)=(x_{1}-x_{2})^{2}+y_{2}^{2}$ is the distance between the two cars, $d_{\text{safe}}$ is the safe distance they aim to maintain, and $\theta_{k}$ is a parameter in the range $[0,100]$ that represents the aggressiveness of each agent (lower $\theta_{k}$ corresponds to more aggressive behavior). Each agent initially assumes that the other agent is a non-aggressive, safety-conscious driver with $\hat{\theta}^{k}_{0}=100$. However, both agents exhibit some level of aggressiveness: the first agent (the blue car) is highly aggressive with $\theta_{1}=5$, while the second agent (the red car) has a moderate aggressiveness level of $\theta_{2}=80$.
To test robustness to misspecified initial beliefs, we introduced an initialization error $e_{init}=\pm 70\%$ in each agent’s guess of the other’s initial estimation and randomly sampled $50$ ”initial parameter estimate of others” for each agent from this error bound. Equivalently, each agent $k$ for each run uniformly samples from the range $\hat{\theta}^{k}\in[30,170]$ and sets it as the initial estimate of agent k-k (although the true initial estimates are $\hat{\theta}^{k}_{0}=100$). Across all runs, as shown in figure 4 a) and b), N-PACE converged to the peer’s true intent and outperformed the baseline, indicating robustness to initialization mismatch. For higher error values, we observed N-PACE starting to show diverging behavior. Also, to better represent the effect of errors in the modeling of peers’ learning dynamics, we used an error $e_{var}=\pm 55\%$ in how each agent models the initial variance in gradient-based Bayesian inference of the other agent as a way to model misspecified convergence speeds of the modeled learning dynamics. We randomly sampled $50$ variances from this range and ran our simulations. Again, across all cases as shown in figure 4 c) and d) we observed N-PACE converging and outperforming the baseline. Indicating how N-PACE can lead to better intent inference despite the mismatch in modeling the initial estimate or learning dynamics of others. Qualitative trajectory comparison results and intent inference Results appear in Fig. 3 and Fig. 4